Problem 9-7 Preferred stock rate of return 58.00, (b) $75.00, (c) What will be the nominal rate of return on a perpetual preferred stock with a $100 par value, a stated dividend of 9% of par, and a current market price of (a) $100.00, and (d) $130.00? Round your answers to two decimal places. a. b. C. d.
Answer #1

Dividend= 100 * .09 = 9

a)Nominal rate of return = dividend / price

                 = 9/ 58

                 = .1552 or 15.52%

b) Nominal rate of return = 9/ 75 = .12 or 12%

c)Nominal rateof return = 9/ 100 = .09 or 9%

d)Nominal rate of return = 9 / 130 = .0692 or 6.92%